# Readability Formulas, Systems and LLMs are Poor Predictors of Reading Ease

### v3

@@ 2.2 Online Reading Measures
We use three primary online measures from the psycholinguistic literature: average Total Fixation Duration (TF), Skip Rate (SR) and Regression Rate (RR). All three measures capture reading ease. Longer reading times and less skipping are associated with increased processing difficulty during reading [47], and with lower language proficiency [12, 1]. Increased regression rates were similarly shown to be a marker of processing difficulty and sentence reanalysis [48, 9, 4]. Correspondingly, more readable texts should have shorter Total Fixation times, more word skipping, and fewer regressions.
@@ 3 Results
Figure 1: Evaluation of readability scoring methods and psycholinguistic measures using their predictivity of reading ease. Years denote the publication year of each method. The evaluations use three eye movement measures of reading ease: TF (the sum of all fixation durations on fixated words), SR (the fraction of skipped words), and RR (the number of backward saccades per word). Each bar depicts the Pearson correlation $r$ coefficient of $\Delta\text{Score}_{M,T}$ with $\Delta\text{ReadingEase}_{E,T}$, where $\text{Score}_{M,T}$ is the readability score difference between pairs of original and simplified texts $T$ according to method $M$, and $\Delta\text{ReadingEase}_{E,T}$ is the average reading ease difference according to measure $E$ on the same pairs of texts across participants. Error bars are 95% confidence intervals. Colors represent the statistical significance level of the correlation.
The results of our main analysis using the OneStopL1&L2 corpus are presented in Figure˜1. Depicted are the Pearson correlations for readability scoring methods (ordered chronologically) and psycholinguistic measures with TF, SR and RR, at the sentence and passage levels. The corresponding pairwise statistical comparisons between the correlation coefficients of all the methods and psycholinguistic measures are presented in Appendix Figure A2. The traditional, modern, and LLM-based readability methods, as well as the two commercial systems, tend to have low and, in many cases, non-significant correlations across textual units and reading ease measures. Notably, we do not observe a pattern of improvement in readability scoring method performance over time. Entropy and the big three word properties are on par or better than readability scoring methods in most evaluations. Their main competitors are Dale-Chall and Coleman-Liau. This is likely related to the strong correlation of these two formulas with word length (see Appendix Figure A3). The advantage of the big three and entropy is especially apparent in the $RR$ evaluation, where the correlations of nearly all other methods are not significant. The best predictor overall tends to be surprisal.
@@ 3.1 Analyses of Generality
Are the surprisal results sensitive to the choice of language model from which the surprisal values were obtained? To answer this question, in Appendix F, we perform the main analysis with 31 additional public language models from 10 language model families. We find that the correlations are stable across models, with very moderate increases as a function of model perplexity for TF and SR and no increase for RR.
@@ 5.5 Eye Movement Measures
    Regression Rate (RR) The number of backward saccades per word.
In Appendix I we provide analyses for additional online measures: First Fixation (FF), Mean Fixation Duration (FD), number of Fixations (NF), first pass measures of Gaze Duration (fpGD), Skip Rate (fpSR) and Regression Rate (fpRR), Gaze Duration (GD), and higher pass Fixation Duration (hpFD). We further include reading speed, an offline measure that can be obtained without eye tracking. Appendix I includes the definitions of all the measures.
@@ Appendix E Analysis of Generality: Results for Different Reading Regimes
In Figures A6 and A7, right panels, we compare the predictivity of surprisal for reading ease using our proposed content-controlled evaluation across different language models, as a function of model log perplexity. We use Pythia 70M from the main analysis, and 31 additional publicly available language models (see list in Table˜A1). At the sentence level, we observe very moderate predictivity increases as function of model log perplexity for TF ($\beta=0.0024$, $p<0.001$) and SR ($\beta=0.0009$, $p<0.001$), and no significant correlation for RR ($\beta=-0.0001$, $p>0.05$), using the linear model $r\sim\text{Perplexity}$. Similar results are obtained at the passage level with TF ($\beta=0.0019$, $p<0.001$), SR ($\beta=0.0024$, $p<0.001$), and RR ($\beta=-0.0009$, $p>0.05$). This suggests that surprisal is a robust predictor of reading ease, irrespective of the language model.
@@ Appendix I Analysis of Generality: Additional Reading Ease Measures
  • First Fixation (FF) The mean duration of the first fixation on a word, considering only words that were not skipped.

### v4

@@ 2.2 Online Reading Measures
We use three primary online measures from the psycholinguistic literature: average Total Fixation Duration (TF), Skip Rate (SR) and Regression Rate ($R$). All three measures capture reading ease. Longer reading times and less skipping are associated with increased processing difficulty during reading [47], and with lower language proficiency [12, 1]. Increased regression rates were similarly shown to be a marker of processing difficulty and sentence reanalysis [48, 9, 4]. Correspondingly, more readable texts should have shorter Total Fixation times, more word skipping, and fewer regressions.
@@ 3 Results
Figure 1: Evaluation of readability scoring methods and psycholinguistic measures using their predictivity of reading ease. Years denote the publication year of each method. The evaluations use three eye movement measures of reading ease: TF (the sum of all fixation durations on fixated words), SR (the fraction of skipped words), and $R$ (the number of backward saccades per word). Each bar depicts the Pearson correlation $r$ coefficient of $\Delta\text{Score}_{M,T}$ with $\Delta\text{ReadingEase}_{E,T}$, where $\text{Score}_{M,T}$ is the readability score difference between pairs of original and simplified texts $T$ according to method $M$, and $\Delta\text{ReadingEase}_{E,T}$ is the average reading ease difference according to measure $E$ on the same pairs of texts across participants. Error bars are 95% confidence intervals. Colors represent the statistical significance level of the correlation.
The results of our main analysis using the OneStopL1&L2 corpus are presented in Figure˜1. Depicted are the Pearson correlations for readability scoring methods (ordered chronologically) and psycholinguistic measures with TF, SR and $R$, at the sentence and passage levels. The corresponding pairwise statistical comparisons between the correlation coefficients of all the methods and psycholinguistic measures are presented in Appendix Figure A2. The traditional, modern, and LLM-based readability methods, as well as the two commercial systems, tend to have low and, in many cases, non-significant correlations across textual units and reading ease measures. Notably, we do not observe a pattern of improvement in readability scoring method performance over time. Entropy and the big three word properties are on par or better than readability scoring methods in most evaluations. Their main competitors are Dale-Chall and Coleman-Liau. This is likely related to the strong correlation of these two formulas with word length (see Appendix Figure A3). The advantage of the big three and entropy is especially apparent in the RRRR evaluation, where the correlations of nearly all other methods are not significant. The best predictor overall tends to be surprisal.
@@ 3.1 Analyses of Generality
Are the surprisal results sensitive to the choice of language model from which the surprisal values were obtained? To answer this question, in Appendix F, we perform the main analysis with 31 additional public language models from 10 language model families. We find that the correlations are stable across models, with very moderate increases as a function of model perplexity for TF and SR and no increase for $R$.
@@ 5.5 Eye Movement Measures
    Regression Rate ($R$) The number of backward saccades per word.
In Appendix I we provide analyses for additional online measures: First Fixation ($F$), Mean Fixation Duration (FD), number of Fixations (NF), first pass measures of Gaze Duration (fpGD), Skip Rate (fpSR) and Regression Rate (fpRR), Gaze Duration (GD), and higher pass Fixation Duration (hpFD). We further include reading speed, an offline measure that can be obtained without eye tracking. Appendix I includes the definitions of all the measures.
@@ Appendix E Analysis of Generality: Results for Different Reading Regimes
In Figures A6 and A7, right panels, we compare the predictivity of surprisal for reading ease using our proposed content-controlled evaluation across different language models, as a function of model log perplexity. We use Pythia 70M from the main analysis, and 31 additional publicly available language models (see list in Table˜A1). At the sentence level, we observe very moderate predictivity increases as function of model log perplexity for TF ($\beta=0.0024$, $p<0.001$) and SR ($\beta=0.0009$, $p<0.001$), and no significant correlation for $R$ ($\beta=-0.0001$, $p>0.05$), using the linear model $r\sim\text{Perplexity}$. Similar results are obtained at the passage level with TF ($\beta=0.0019$, $p<0.001$), SR ($\beta=0.0024$, $p<0.001$), and $R$ ($\beta=-0.0009$, $p>0.05$). This suggests that surprisal is a robust predictor of reading ease, irrespective of the language model.
@@ Appendix I Analysis of Generality: Additional Reading Ease Measures
  • First Fixation ($F$) The mean duration of the first fixation on a word, considering only words that were not skipped.